A metal alloy weighing 9 lb. and containing 75% copper is melted and mixed with 11 lb. of a different alloy which contains 15% copper. What percent of the resulting alloy is copper?
Answer by stanbon(75887)

A metal alloy weighing 9 lb. and containing 75% copper is melted and mixed with 11 lb. of a different alloy which contains 15% copper. What percent of the resulting alloy is copper?
-------------------
Equation:
copper + copper = copper
0.75*9 + 0.15*11 = x*20 lb
------
75*9 + 15*11 = x*2000
-----
x = (675 + 161)/2000
x = 41.8% (percent of copper in the final mix)
---------------
Cheers,
Stan H.
-----------------